OpenStudy (anonymous):

Solve the following system of equations. Enter the y-coordinate of the solution. Round your answer to the nearest tenth. 5x+2y=21 -2x+6y=-34

OpenStudy (anonymous):

lets try substitution 5x+2y=21 =5x=21-2y =x=(21-2y)/5 sub into second equation

OpenStudy (anonymous):

-2(21-2y)/5+6y=-34 (-42+4y)/5+6y=-34 solve for y y=-64/17 plug into equation and solve for x=(97/17) weird.

OpenStudy (anonymous):

theres your answer. checked it. it's correct. weird retricenumbers but correct

OpenStudy (anonymous):

hm when i did it i did -2(21-2y)/5+6y=-34 solves for y + -41+4y/5+6y=-34 the get rid of the fraction so i multiplied both sides by 5/1 and i got -41+4y+64=170 then combine like terms and add 41 to both sides and i got y=21.1

OpenStudy (anonymous):

OH WAIT i just saw my mistake

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

but im still a lil confused how did you get -64/17

OpenStudy (anonymous):

like what method did you use to solve for x

OpenStudy (anonymous):

i mean y

OpenStudy (anonymous):

i did subtitution for y and then plugged y into any of the other equations to get x

OpenStudy (anonymous):

like why are our answers so different bc i redid it and i got 21.2

OpenStudy (anonymous):

clear fractions (-42+4y)+30y=-170 =34y=-170+42 =34y=-128 =y=-64/17

OpenStudy (anonymous):

does 21.2 work when you plug it in?

OpenStudy (anonymous):

y=-64/17=-3.76 x=97/17=5.71

OpenStudy (anonymous):

It's -3.8